### v3

@@ 3.2 OR-Instruct: Towards Training Effective OR LLMs
Initially, OR-Instruct attempts to generate new data by expanding scenarios and question types from the training data pool in a bootstrapping fashion with GPT-4. In our particular case, we start with 686 real-world industry cases collected from some OR textbooks and our previous industrial projects (after proper abstraction and anonymization). For each generation, we sample 3 entries from this pool as in-context examples, keeping the input token length suitable for GPT-4. Of the 3 entries, 2 are from real-world entries and 1 is from the model-generated entries in previous iterations, if available, to promote diversity. The prompting template is shown in Appendix 8.3.
This approach is similar to general synthesis methods such as Wang et al. (2022), but OR-Instruct focuses on expanding scenarios rather than tasks. We observe that this step partially addresses the Comprehensive Coverage requirement in terms of scenarios and question types. However, it falls short of meeting the other requirements, especially concerning varying levels of difficulty. In manually reviewing the difficulty of generated example, of 50 cases, 87% are deemed easy, 13% medium, and none hard, as judged by criteria in Appendix 8.1. We also provide examples in Appendix 8.9 for comparing generated easy entries with real-world hard entries. Fortunately, the original seed data already shows a diverse range of difficulties. Thus, we can naturally enhance the difficulty diversity by augmenting them described in the next strategy, thereby further addressing Comprehensive Coverage.
@@
    Altering Objectives and Constraints: The first augmentation involves adding, removing, or replacing objectives and constraints in the problem, along with making necessary adjustments to the mathematical models and programs. Specifically, we start by providing GPT-4 with the original example and ask it to list five potential changes to the objectives and constraints. These suggested changes are then fed back to GPT-4 using a prepared few-shot prompt to modify the problem, model, and programs accordingly. This augmentation is designed to enhance Environmental Adaptability. An example is provided in Figure 4, and the corresponding prompting template is shown in Appendix 8.4.
@@
    Rephrasing Questions: The second augmentation modifies the description of the question, simulating the expression habits of different users. This process involves instructing GPT-4 to rewrite the target problem, either simplifying or complicating it, while ensuring the core logic aligns with the solution, including the mathematical model and programs. This augmentation is designed to enhance Linguistic Diversity. An example is provided in Figure 5, and the corresponding prompting template is shown in Appendix 8.5.
@@
    Incorporating Multiple Modeling Techniques: The third augmentation explores the use of different modeling techniques. We identify five potential techniques from engineers’ experiences, such as introducing auxiliary variables or using the Big $M$ method, for GPT-4 to choose from in modifying an objective or constraint in the original mathematical model. This augmentation is designed to enhance Solution Variability. An example is provided in Figure 6, and the corresponding prompting template is shown in Appendix 8.6.
@@
Finally, OR-Instruct implements several heuristic measures on the generated examples. Initially, we eliminate examples that contain duplicate questions or those that overlap with questions from our evaluation benchmarks. Subsequently, we apply a regular match correction function to amend minor grammatical errors in the programs, which are occasionally induced by GPT-4’s limited familiarity with the COPT API (refer to Appendix 8.2). Training data that contain unexecutable programs are discarded, as they represent manifestly low-quality data. After conducting a manual review of the remaining data, which demonstrates an acceptable accuracy rate — 70% for expansion data and 75% for augmentation data — we opt to bypass further filtering in pursuit of establishing a fully automatic filtering mechanism. This final filtering step is projected to exclude approximately 39% of the generated examples, with the surviving examples subsequently incorporated into our training data pool.
@@ 4 Results
In this section, we train the open-source 7b size large language models with synthetic data from OR-Instruct and compare the results with GPT-4 as well as benchmarks. Then, using the case of mixed integer linear programming as an example, we demonstrate the customized enhancement effect of OR-Instruct by including directionally generated training data.
@@ 4.1 Data Generation by Using OR-Instruct
4.2 Model Training and Inference
We borrow a widely used training framework from open-instruct (Wang et al. 2023d.) where we take a natural language OR problem wrapped in an Alpaca-like template (Taori et al. 2023.) as the input prompt (see Appendix 8.7), and treat a complete solution including mathematical models and programs as the target completion. During training, we compute the loss only over the target completion. We apply OR-Instruct Data to train several open-source LLMs of approximately 7b size, including Mistral-7B (Jiang et al. 2023.), Deepseek-Math-7B-Base (Shao et al. 2024.), and LLaMA-3-8B (AI@Meta 2024.), as the backbones. We list the hyper-parameters for each backbone in Table 1. We call the resulting models as ORLMs. For inference during evaluation, we always use greedy decoding under 0-shot setting to eliminate randomness, selecting the top-1 completion as the final solution. We extract the program from the solution and execute it to obtain the predicted optimal value.
@@ 4.3 Evaluation and Baselines
Evaluation Benchmarks and Metrics: We use NL4OPT (Ramamonjison et al. 2023.), MAMO (Huang et al. 2024b.), and IndustryOR as evaluation benchmarks. NL4OPT is the most widely used benchmark for operations research and includes 289 linear programming problems in its test set. However, NL4OPT only provides mathematical models as targets, which complicates the verification of execution accuracy due to the absence of optimal solutions. To address this, we convert these mathematical models into programs using GPT-4, calculate and check the optimal solutions, and use these as ground truth. MAMO, a concurrent project, evaluates the mathematical modeling capabilities in LLMs. It includes 652 easy and 211 complex linear programming problems, each paired with its corresponding optimal solution, sourced from various academic materials. IndustryOR, which is proposed in this work, is the first industrial benchmark, consisting of 100 real-world OR problems from eight industries. It covers 5 types of questions linear programming, integer programming, mixed integer programming, non-linear programming, and others across 3 levels of difficulty. We measure performance using execution accuracy, where an executed optimal value that matches any provided ground truth optimal value is considered correct. Compared to NL4OPT, this metric enables a fully automated evaluation and provides greater flexibility for mathematical modeling approaches.
@@
Standard 42.4%* - - - - -
Reflexion 50.7%* - - - - -
Chain-of-Experts 58.9%* - - - - -
@@
Standard 47.3%* 66.5%* 14.6%* 28.0% 50.2% 39.1%
Reflexion 53.0%* - - - - -
Chain-of-Experts 64.2%* - - - - -
OptiMUS 78.8%* - - - - -
@@
ORLM-Mistral-7B 84.4% 81.4% 32.0% 27.0% 68.8% 56.2%
ORLM-Deepseek-Math-7B-Base 86.5% 82.2% 37.9% 33.0% 71.2% 59.9%
ORLM-LLaMA-3-8B 85.7% 82.3% 37.4% 38.0% 71.4% 60.8%
The results are presented in Table 2. First, it is clear that methods based on LLMs generally outperform the PLM-based best method (tag-BART) in the NL4OPT test. This suggests that PLMs have limited generalization capabilities. For proprietary LLMs, as the mathematical reasoning capability increases from GPT-3.5 to GPT-4, we observe that performance has advanced across all prompt engineering methods. Finally, ORLMs based on various open-source LLMs have demonstrated significantly improved optimization modeling capabilities, compared to vanilla open-source LLMs, which score 0 for all benchmarks in our preliminary experiments due to failing to output executable programs. This underscores the effectiveness of the OR-Instruct data. Our best-performing ORLM, trained on LLaMA-3-8B, achieves state-of-the-art performance on NL4OPT, MAMO (including both easy and complex linear programming), and IndustryOR benchmarks, surpassing Standard prompting based on GPT-4 by 42.2% in micro average [^1] Micro average aggregates the outcomes of all classes to compute the metrics, heavily weighing the classes with more instances. This method is effective for evaluating the overall effectiveness of the model, particularly in datasets where some classes significantly outnumber others. and 55.4% in macro average [^2] Macro average calculates individual metrics for each class without regard to class frequency, and then averages these metrics, treating all classes with equal importance. This approach highlights model performance on minority classes and is beneficial for assessing model fairness across diverse class distributions.. We also want to highlight that, unlike tag-BART, we ensure no training and validation data from NL4OPT was used. Thus, the performances of the ORLMs across all benchmarks reflect out-of-domain generalization.
@@ 4.4 Customization Enhancement
In this subsection, we illustrate the capability of our framework to enhance large models in specific areas by generating targeted types of optimization modeling data. Specifically, we focus on the scenario of mixed integer linear programming (MILP). We provide an empirical example demonstrating how the quality of responses improves with training. Initially, we use a base dataset comprising 10,000 general modeling cases. To enhance the ORLM’s proficiency in MILP, we selected 10 fundamental MILP questions and their answers for inclusion in the seed data pool. After several iterative rounds, this process yields 2,000 tailored training data points for MILP. We will subsequently assess the impact of incorporating these 2,000 data points into the training dataset by comparing the ORLM’s performance before and after their addition.
@@ 5.1 Detailed Comparison of ORLM vs GPT-4 on IndustryOR
To assess the optimization modeling capabilities across different levels of difficulty and question types, we compare ORLM-LLaMA-3-8B and Standard-GPT-4 on IndustryOR as shown in Figure 9. ORLM-LLaMA-3-8B shows superior performance over Standard-GPT-4 across all difficulty levels, especially in the hard category. Regarding different question types, ORLM-LLaMA-3-8B outperforms Standard-GPT-4 in linear programming, integer programming, and mixed-integer programming. Both models perform poorly in non-linear programming and other rare question types. This is partly due to the fact that the seed dataset collected contains too few examples of nonlinear and other types of optimization modeling, resulting in a low percentage of that type in the training dataset, and partly due to the fact that the inherent complexity and scarcity of these types of questions. Overall, the OR-Instruct data proves effective in enhancing Comprehensive Coverage across various question types and difficulty levels.
Figure 9: Comparison of ORLM and GPT-4 on IndustryOR across different difficulty levels and question types.
5.2 Ablation Study on OR-Instruct Augmentations
To further verify the effectiveness of the augmentations in OR-Instruct, we conduct detailed ablation experiments to study the effects of altering objectives and constraints, rephrasing questions, and incorporating multiple modeling techniques. Specifically, we select a controlled sample size of 3,000 with certain augmentations from the OR-Instruct data, and then use these to train the LLaMA-3-8B model with the same hyper-parameters across all experiments. The results are presented in Table 3. Training data with all three augmentations (denoted as Full Augmentations) achieves a base performance of 68.6% in micro average and 55.7% in macro average. Removing any of the three augmentations from the base setting leads to a performance drop across all benchmarks, both in micro and macro averages. Rephrasing questions seems slightly more important than the other two. Overall, the results show that all three augmentations contribute to general performance, proving their effectiveness in enhancing the data diversity.

### v4

@@ 3.2 OR-Instruct: Towards Training Effective OR LLMs
Initial Dataset
The quality of initial data plays a crucial role in large model training, providing a strong foundation for diversity and difficulty as a starting point (Marion et al. 2023). Additionally, in this customizable framework designed to enhance modeling capabilities, the category or domain of the optimization problem associated with the initial data is important. We will demonstrate later that when the initial dataset is concentrated in a specific type (e.g., scenario or type), it substantially enhances the LLM’s ability to model effectively within that domain.
In our particular case, we start with 686 real-world industry cases collected from some OR textbooks and our previous industrial projects (after proper abstraction and anonymization).
@@
Initially, OR-Instruct seeks to generate new data by expanding scenarios and question types from the training dataset using a bootstrapping approach powered by GPT-4. The process begins with GPT-4 generating a list of 100 scenarios where optimization models are applied in real-life contexts. In each iteration, three examples are selected from data pool to serve as in-context references for GPT-4, and one scenario is randomly chosen from the list. Among the three examples, two are sourced from real-world data, while the third, if available, is taken from data previously generated by the model. Finally, GPT-4 is prompted to create a new sample based on the three examples within the context of the selected scenario. This iterative approach promotes greater diversity in the resulting dataset. The prompting template for this expansion process is detailed in Appendix 8.2.
It’s important to note that optimization models are naturally abstract, often built on similar logical ideas across various scenarios. In this setting, Comprehensive Coverage not only exposes the LLMs to a broad range of situations but also helps it better grasp the underlying patterns and principles when modeling across these different contexts. However, it falls short of meeting the other requirements, especially concerning varying levels of difficulty. In manually reviewing the difficulty of generated example, of 50 cases, 87% are deemed easy, 13% medium, and none hard, as judged by criteria in Appendix 8.1. We also provide examples in Appendix 8.6 for comparing generated easy entries with real-world hard entries. Fortunately, the original seed data already shows a diverse range of difficulties. Thus, we can naturally enhance the difficulty diversity by augmenting them described in the next strategy, thereby further addressing Comprehensive Coverage.
@@
    Altering Objectives and Constraints: The first augmentation involves adding, removing, or replacing objectives and constraints in the problem, along with making necessary adjustments to the mathematical models and programs. Specifically, we start by providing GPT-4 with the original example and ask it to list five potential changes to the objectives and constraints. These suggested changes are then fed back to GPT-4 using a prepared few-shot prompt to modify the problem, model, and programs accordingly. This augmentation is designed to enhance Environmental Adaptability. An example is provided in Figure 4, and the corresponding prompting template is shown in Appendix 8.3.1.
    Remark 3.1
    One important point to note is that during the process of altering objectives and constraints, GPT-4 may generate background scenarios that are illogical, such as presenting two contradictory constraints. However, this issue arises from unreasonable problem descriptions. As long as the model can accurately map each statement of the problem to the correct objectives and constraints, it can still yield the desired results when tested on meticulously designed datasets that reflect such issues.
@@
    Rephrasing Questions: The second augmentation involves modifying the problem description while essentially addressing the same optimization model. This approach enhances the robustness of large language models (LLMs) to different prompts, enabling them to abstract a unified optimization model across various problem descriptions. This process involves instructing GPT-4 to rewrite the target problem, either simplifying or complicating it, while ensuring the core logic aligns with the solution, including the mathematical model and programs. This augmentation is designed to enhance Linguistic Diversity. An example is provided in Figure 5, and the corresponding prompting template is shown in Appendix 8.3.2.
@@
    Incorporating Multiple Modeling Techniques: The third augmentation explores the use of different modeling techniques. We identify five potential techniques from engineers’ experiences, such as introducing auxiliary variables or using the Big $M$ method, for GPT-4 to choose from in modifying an objective or constraint in the original mathematical model. This augmentation is designed to enhance Solution Variability. An example is provided in Figure 6, and the corresponding prompting template is shown in Appendix 8.3.3.
@@
At the end of each iteration, OR-Instruct implements correction and filtering measures on the generated examples. First, we apply a regular match correction function to address minor grammatical errors in the programs, which may arise from GPT-4’s limited familiarity with the COPT API. Unexecutable programs are discarded, as they clearly represent low-quality data. Notably, in the final iteration, we also eliminate duplicate questions and any examples overlapping with the evaluation benchmarks to ensure the dataset remains uncontaminated.
Based on manual sampling assessments, we find that the accuracy of the synthesized data by expansion operation is about 70% and 75% for the augmentation data, judged by the correctness of both the code and the model. On average, each iteration filters out approximately 39% of the generated examples, with the remaining examples incorporated into our training data pool.
In summary, the above outlines all the key steps of OR-Instruct, and the specific procedures can be found in Figure 2.
@@ 4 Results
In this section, we train the open-source 7b size large language models with synthetic data from OR-Instruct and compare the results with GPT-4 as well as other benchmarks. Then, using the case of mixed integer linear programming as an example, we demonstrate the customized enhancement effect of OR-Instruct by including directionally generated training data.
@@ 4.1 Data Generation by Using OR-Instruct
4.2 Model Finetuning and Inference
This paper employs instruction tuning for open-source large models, a technique designed to enhance the ability of large language models (LLMs) to follow diverse natural language instructions. The primary aim is to improve zero-shot performance on unseen tasks, facilitating effective generalization across various instruction-based scenarios. The methodology involves training on a dataset of instruction-output pairs, typically consisting of three elements: an instruction (e.g., "Translate this sentence from English to Spanish"), optional context, and the target output. The training can be mathematically represented as:
$\mathcal{L}(\theta)=-\sum_{(x,y)\in\mathcal{D}}\log P(y\mid x;\theta)$
where $\mathcal{L}(\theta)$ is the loss function, $x$ is the instruction input, $y$ is the expected output, and $\theta$ denotes the model parameters. We borrow a widely used training framework from open-instruct (Wang et al. 2023d), where we take a natural language OR problem wrapped in an Alpaca-like template (Taori et al. 2023.) as the input prompt (see Appendix 8.4), and treat a complete solution including mathematical models and programs as the target completion.
During training, we compute the loss only over the target completion. We employ the OR-Instruct dataset to train several open-source large language models (LLMs) of approximately 7 billion parameters, including Mistral-7B (Jiang et al. 2023), Deepseek-Math-7B-Base (Shao et al. 2024), and LLaMA-3-8B (AI@Meta 2024), which serve as our backbone models. We conduct a grid hyper-parameter search for each backbone and present the best-performing hyperparameters in Table 1. Prior to the final fine-tuning, we conducted a grid search with learning rates within {2e-5, 5e-5, 7e-5, 3e-6, 5e-6, 7e-6}, batch sizes within {64, 128, 256, 512}, and training epochs within {1, 2, 3} to determine the optimal settings. The models resulting from this process are referred to as ORLMs.
Regarding the inference techniques during evaluation, it is important to consider that large language models, as probabilistic generative models, exhibit inherent randomness. This randomness can lead to inconsistent outcomes across multiple evaluations. To mitigate the influence of this randomness, we implemented a greedy decoding strategy in a zero-shot scenario, consistently selecting the highest probability output (top-1) as the final result. After obtaining these results, we executed the corresponding code to derive the predicted optimal values for comparison with the ground truth. Although this greedy approach may reduce the performance of large models on such tasks, it ensures the stability of the output and enables the reproducibility of our findings under identical conditions. We will discuss the impact of employing additional inference techniques on the performance of large models in Section 5.4.
@@ 4.3 Evaluation and Baselines
Evaluation Benchmarks and Metrics: We use NL4OPT (Ramamonjison et al. 2023.), MAMO (Huang et al. 2024b.), and IndustryOR as evaluation benchmarks. NL4OPT is the most widely used benchmark for operations research and includes 289 linear programming problems in its test set. However, NL4OPT only provides mathematical models as targets, which complicates the verification of execution accuracy due to the absence of optimal solutions. To address this, we convert these mathematical models into programs using GPT-4, calculate and check the optimal solutions, and use these as ground truth. MAMO, a concurrent project, evaluates the mathematical modeling capabilities in LLMs. It includes 652 easy and 211 complex linear programming problems, each paired with its corresponding optimal solution, sourced from various academic materials. IndustryOR, which is proposed in this work, is the first industrial benchmark, consisting of 100 real-world OR problems from eight industries. It covers 5 types of questions, linear programming, integer programming, mixed integer programming, non-linear programming, and others across 3 levels of difficulty. We measure performance using execution accuracy, where an executed optimal value that matches any provided ground truth optimal value is considered correct. Compared to NL4OPT, this metric enables a fully automated evaluation and provides greater flexibility for mathematical modeling approaches.
@@
In addition, we introduce the most well-known open-source large models for comparison, including Llama-3.1-lnstruct (Dubey et al. 2024), DeepSeek-v2-Chat (DeepSeek AI 2024), Qwen2-Instruct (Qwen Team 2024), and Mistral-Nemo (Mistral AI 2024). Each model is tasked with constructing optimization models and generating code to solve the test set problems.
We also assess the difference between the performance of ORLM and human-level performance. Given that constructing and solving optimization models requires a solid foundation in operations research and programming skills, we recruit 8 highly capable senior undergraduate students from relevant fields and 8 experts in the domain to evaluate four types of test sets. Each test set is evaluated by 2 senior undergraduate students and 2 experts. Each participant randomly selects 70 questions from their corresponding test set and is required to construct an optimized model and calculate the optimal solution. Participants are allowed to use any programming language; however, to ensure fairness in comparison with large models, they are only permitted to submit a finalized version of their code for testing and cannot make further modifications. Finally, the accuracy rates for the student and expert groups on each test set are obtained by averaging the results.
@@
Method/Model Size NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
@@
Standard Unknown 42.4%* - - - - -
Reflexion Unknown 50.7%* - - - - -
Chain-of-Experts Unknown 58.9%* - - - - -
@@
Standard Unknown 47.3%* 66.5%* 14.6%* 28.0% 50.2% 39.1%
Reflexion Unknown 53.0%* - - - - -
Chain-of-Experts Unknown 64.2%* - - - - -
OptiMUS Unknown 78.8%* - - - - -
Standard prompting based on open-source LLMs
Llama-3.1-Instruct 405B 38.7% 35.1% 20.8% 13.0% 31.5% 26.9%
DeepSeek-V2-Chat 236B 66.5% 60.5% 32.7% 16.0% 53.1% 43.9%
Qwen2-Instruct 72B 72.6% 79.9% 29.0% 18.0% 64.4% 49.8%
Mistral-Nemo 12B 14.6% 19.4% 3.7% 7.0% 14.6% 11.1%
@@
ORLM-Mistral 7B 84.4% 81.4% 32.0% 27.0% 68.8% 56.2%
ORLM-Deepseek-Math 7B 86.5% 82.2% 37.9% 33.0% 71.2% 59.9%
ORLM-LLaMA-3 8B 85.7% 82.3% 37.4% 38.0% 71.4% 60.8%
ORLM-Qwen2.5 7B 86.1% 85.2% 44.1% 25% 73.7% 60.1%
Human Evaluation
Senior Undergraduates - 80.4% 84.9% 53.1% 44.0% 75.2% 65.6%
Experts - 94.3% 90.4% 78.9% 76.0% 85.0% 88.2%
The results are presented in Table 2. First, it is clear that methods based on LLMs generally outperform the PLM-based best method (tag-BART) in the NL4OPT test. This suggests that PLMs have limited generalization capabilities. For proprietary LLMs, as the mathematical reasoning capability increases from GPT-3.5 to GPT-4, we observe that performance has advanced across all prompt engineering methods.
Our best-performing ORLMs, like ORLM-LLaMA-3 and ORLM-Qwen2.5, achieve state-of-the-art performance across four benchmarks with respect to GPT-4 and other open-source LLMs, surpassing Standard prompting based on GPT-4 by 42.2% in micro averageMicro average aggregates the outcomes of all classes to compute the metrics, heavily weighing the classes with more instances. This method is effective for evaluating the overall effectiveness of the model, particularly in datasets where some classes significantly outnumber others. and 55.4% in macro averageMacro average calculates individual metrics for each class without regard to class frequency, and then averages these metrics, treating all classes with equal importance. This approach highlights model performance on minority classes and is beneficial for assessing model fairness across diverse class distributions.. A comparison with human evaluation results reveals that ORLMs outperform senior undergraduate students on simpler problems (NL4OPT/MAMO-EasyLP) and approach expert-level performance. However, for more complex problems (IndustryOR/MAMO-ComplexLP), ORLMs still underperform relative to the average student, indicating that their problem-solving capabilities for complex tasks remain inadequate. We will analyze the underlying reasons for this in Section 5, highlighting that ORLMs still have the potential to surpass student performance and approach expert-level solutions on more challenging problems.
While ORLMs demonstrate strong performance in optimization modeling tasks, we also evaluate their efficacy on general large language model benchmarks after domain-specific fine-tuning. In particular, we conduct a rigorous evaluation of the LLaMA-3-8B model (pre-fine-tuning) and the ORLM-LLaMA-3-8B model (post-fine-tuning) following the open-instruct (Wang et al. 2023d) on benchmarks from various domains, including mathematics (GSM8K), code (HumanEval), and general knowledge (MMLU, BBH, TydiQA), the results are listed in Table 3.
Table 3: Comparison of accuracy rate between LLaMA-3-8B and ORLM-LLaMA-3-8B across multiple benchmark domains.
Model GSM8K HumanEval MMLU BBH TydiQA
LLaMA-3-8B 56.5% 67.7% 65.2% 63.7% 21.1%
ORLM-LLaMA-3-8B 58.0% 70.6% 64.6% 61.9% 21.9%
The results reveal that ORLM-LLaMA-3-8B exhibits improved performance in the mathematics and coding domains, which aligns with expectations, as the OR-Instruct data emphasizes mathematical modeling and coding with solver assistance. However, the model exhibits a slight decline in general domain performance. Overall, the average performance across all domains shows no substantial deviation.
@@ 4.4 Customization Enhancement
In this subsection, we illustrate the capability of our framework to enhance large models in specific areas by generating targeted types of optimization modeling data. Specifically, we focus on the scenario of mixed integer linear programming (MILP). We provide an empirical example demonstrating how the quality of responses improves with training. Initially, we use a base dataset comprising 10,000 general modeling cases. To enhance the ORLM’s proficiency in MILP, we selected 50 fundamental MILP questions and their answers for inclusion in the seed data pool. After several iterative rounds, this process yields 2,000 tailored training data points for MILP. We will subsequently assess the impact of incorporating these 2,000 data points into the training dataset by comparing the ORLM’s performance before and after their addition.
@@ 5.1 Detailed Comparison of ORLM vs GPT-4 on IndustryOR
To assess the optimization modeling capabilities across different levels of difficulty and question types, we compare the accuracy rate of ORLM-LLaMA-3-8B and Standard-GPT-4 on IndustryOR, as shown in Figure 9. ORLM-LLaMA-3-8B shows superior performance over Standard-GPT-4 across all difficulty levels, especially in the hard category. Regarding different question types, ORLM-LLaMA-3-8B outperforms Standard-GPT-4 in linear programming, integer programming, and mixed-integer programming. Both models perform poorly in non-linear programming and other rare question types. This is partly due to the fact that the seed dataset collected contains too few examples of nonlinear and other types of optimization modeling, resulting in a low percentage of that type in the training dataset, and partly due to the fact that the inherent complexity and scarcity of these types of questions. Overall, the OR-Instruct data proves effective in enhancing Comprehensive Coverage across various question types and difficulty levels.
Figure 9: Accuracy rate of ORLM and GPT-4 on IndustryOR across different difficulty levels and question types.
5.2 Ablation Study
Ablation Study on Data Synthesis Strategy
We first assess whether the data generated by OR-Instruct truly leads to improvements in the Seed data, particularly in terms of the performance of the ORLM. Particularly, we compare the performance of fine-tuning on seed data alone with the results of fine-tuning on OR-Instruct-generated data based on the same seed data. We observe that, with the exception of minor variations in simpler linear programming cases (MAMO_EasyLP), the OR-Instruct-generated data yield a significant performance boost for ORLM than only using the seed data as shown in Table 4.
Table 4: Performance comparison with varying data sizes
Base Model Data Size NL4OPT MAMO EasyLP MAMO ComplexLP IndustryOR Micro Avg Macro Avg
LLaMA-3-8B 32,481 (Full data) 85.7% 82.3% 37.4% 38.0% 71.4% 60.8%
LLaMA-3-8B 686 (seed data) 75.1% 82.8% 29.4% 20.0% 67.0% 51.8%
Ablation Study on OR-Instruct Augmentations
To further verify the effectiveness of the augmentations in OR-Instruct, we conduct detailed ablation experiments to study the effects of altering objectives and constraints, rephrasing questions, and incorporating multiple modeling techniques. Specifically, we construct four distinct datasets, each comprising 3,000 instances, drawn from the OR-Instruct data but employing different augmentation strategies. These datasets are then utilized to train the LLaMA-3-8B model, maintaining consistent hyperparameters across all experimental conditions. The results are presented in Table 5. Training data with all three augmentations (denoted as Full Augmentations) achieves a base performance of 68.6% in micro average and 55.7% in macro average. Removing any of the three augmentations from the base setting leads to a performance drop across all benchmarks, both in micro and macro averages. Rephrasing questions seems slightly more important than the other two. Overall, the results show that all three augmentations contribute to general performance, proving their effectiveness in enhancing the data diversity.